1. (2) At pH 5.0, what is the concentration of OH- in solution? ___________

2. (2) If the monoprotic acid HAb has a pK of 3 and a solution is adjusted to pH 5:

a) what is the concentration of hydrogen ions in that solution? ____________

b) what is the ratio of Ab- to HAb? ________________

3. (2) Oxalic acid, a di-protic, di-carboxylic acid has pKs of 1.3 and 4.3 and thus can exist as H2Ox, HOx- and Ox= (please use the same in your answer) .

a) What would be the dominant form at pH 3.3? ________________

            b) What would be the second most dominant from at pH 3.3? ___________
Answer #1

1)At pH = 5, the hydrogen ion concentration is [H+] = 1 x 10-5M thus [OH-] = Kw / [H+]

         [OH-] = 1.0x 10-14 / 1.0 x 10-5 = 1.0 x 10 -9 M

2) pKa of the3 acid = 3.0 which shows it is relatively weak acid.

   HA <---> H+ +A- and 1.0x 10-3 = Ka = [H+] [A-] / [HA]

a) When the pH = 5.0 , the [H+} = 1.0x 10-5 M as the pH is adjusted to pH= 5.0

b) the ratio [A-] / [HA] = Ka/ [H+] = {1.0x 10 -3] / 1.0 x 10-5 = 100

   Since the pH is increased to 5.0 which is greater than its pKa, the acid is more dissociated.

3) The acid is a dicarboxylic acid with pK values, 1.3 and 4.3.

   When the aqueous solution of H2Ox has pH = Pk1 = 1.3 , exactly 50% of one -H is replaced that is we have 50% H2Ox and 50% HOx- . similarly at pH = Pk2 = 4.3 we have 50% of HOx- and 50% of Ox=.

Thus at a mid pH of 2.8 = (1.3+4.3)/2 , we have 100% HOx-.

a) The dominant form at pH = 3.3 will be HOx- [since the pH crossed the 100% HOx- pH but not reached 50% Ox-=]

b) the second most dominat form is Ox= since the pH is nearing the pK2.